What is the circumference of a circle with a radius of 4.1 m? Round to the nearest tenth.Use pi=3.14




Thanks,
Amanda
: What is the circumference of a circle with a radius of 4.1 m? Round to the nearest tenth.Use pi=3.14




Thanks,
Amanda

Answer by stanbon(19020)

What is the circumference of a circle with a radius of 4.1 m? Round to the nearest tenth.Use pi=3.14
-----------------
Circumference = 2(pi)r
C = 2(3.14)4.1 = 25.748
===================
Cheers,
Stan H.
Question 149615: What is the circumference of a circle with a radius of 4.1 m? Round to the nearest tenth.Use pi=3.14




Thanks,
Amanda
: What is the circumference of a circle with a radius of 4.1 m? Round to the nearest tenth.Use pi=3.14




Thanks,
Amanda

Answer by jim_thompson5910(9401)


C=2*pi*r Start with the circumference of a circle formula


C=2*pi*4.1 Plug in r=4.1


C=25.76106 Multiply 2, pi and 4.1 (you can use 3.14 for pi) to get 25.76106



So the circumference of a circle with a radius of 4.1 meters is 25.8 meters (rounded to the nearest tenth)